OpenStudy (anonymous):

solve the following equation 3x+2y=18 y=-2/3x+12

OpenStudy (anonymous):

Hi ! Have you a way ?! If you have please write it ;)

OpenStudy (anonymous):

There are multiple ways of solving this set of equations. Have you tried to do any? challenging

OpenStudy (anonymous):

have answer but cannot figure how they got it

OpenStudy (anonymous):

I have solved tons of these so I know a few tricks. 3x+2y=18 y=-2/3x +12 You can simply write 3x+2*(-2/3x+12)=18 as the second equation already had y expressed using only x

OpenStudy (anonymous):

Now this equation has only x and real numbers so you should be able to express x. Once thats done using the y= equation you can find y as well

OpenStudy (anonymous):

3x+2y=18 2y=18-3x y=9-3/2x

OpenStudy (texaschic101):

y = -2/3x + 12 now sub -2/3x + 12 in for y in the other equation 3x + 2y = 18 3x + 2(-2/3x + 12) = 18 -- distribute through the parenthesis 3x - 4/3x + 24 = 18 -- subtract 24 from both sides 3x - 4/3x = 18 - 24 -- combine like terms (common denominator is 3) 9/3x - 4/3x = - 6 5/3x = -6 -- divide by 5/3 x = -6/(5/3) x = -6 * 3/5 x = - 18/5 now sub -18/5 in for x in either of the original equations y = -2/3x + 12 y = -2/3(-18/5) + 12 y = 36/15 + 12 -- common denominator is 15 y = 36/15 + 180/15 y = 216/15 check... 3x + 2y = 18 3(-18/5) + 2(216/15) = 18 - 54/5 + 432/15 = 18 -- common denominator is 15 -162/15 + 432/15 = 18 270/15 = 18 18 = 18 (correct) x = -18/5 and y = 216/15

OpenStudy (anonymous):